# Measuring all the noises of LLM Evals

### v1

@@ Abstract
Separating signal from noise is central to experimental science. Applying well-established statistical method effectively to LLM evals requires consideration of their unique noise characteristics. We clearly define and measure three types of noise: prediction noise from generating different answers on a given question, data noise from sampling questions, and their combined total noise following the law of total variance. To emphasize relative comparisons and gain statistical power, we propose the all-pairs paired method, which applies the paired analysis to all pairs of LLMs and measures all the noise components based on millions of question-level predictions across many evals and settings. These measurements revealed clear patterns. First, each eval exhibits a characteristic and highly predictable total noise level across all model pairs. Second, paired prediction noise typically exceeds paired data noise, which means reducing prediction noise by averaging can significantly increase statistical power. These findings enable practitioners to assess significance without custom testing and to detect much smaller effects in controlled experiments.
@@ 1 Introduction
Separating signal from noise is central to experimental science. Well-established methods from statistics can increase the rigor of the conclusions drawn from data benjamini2021asa. Even though these methods are in landmark textbooks and clearly described for LLM evals miller2024error, applying them effectively still requires some consideration of the prediction noise uniquely important for rich generative models like LLMs. Otherwise much potential power is lost (Example 1) and confusions ensue. Unlike previous probabilistic models that typically have a clear best prediction, LLMs produce persistently diverse predictions. Unlike physical experiments where treatments cannot be undone, digital LLMs allow for repeated independent samples, thus allowing us to directly measure their prediction noise. To better understand the sources of noise, we decompose the total variance into two interpretable components using the law of total variance. For a model $A$ evaluated on questions $x$ with prediction seed $\epsilon$, we have
@@
The prediction variance $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ is due to the model generating different answers on a fixed question. The data variance $\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A]]$ measures how the questions vary in difficulty level and thus vary in their expected metrics. Prediction variance can be reduced but usually remains persistent; Data variance matters for reliability beyond the specific question set; The total variance is what matters in an experiment and is available even without multiple metrics per question.
While prediction variance can be reduced by averaging, the data variance is irreducible and can be large when the questions have different difficult levels. However, similar models may have similar performance on each question, suggesting that greatly reduced data noise is only possible with the paired analysis. A difficulty is that the paired method only applies to pairs of models rather than a single model or a list of models. In this work, we propose the all-pairs paired method to measure all the noises, which use the paired analysis on all pairs of LLMs to measure the data, prediction, and total paired variances $\mathrm{Var}[A-B]$, for all model pairs $A,B$, and across many evals and settings, revealing clear patterns.
@@
To be more accurate, Figure 1 Left shows the paired total noise on SWEBench-Verified against our theoretical prediction. This allows us to have accurate-enough noise measurements without custom analysis and even to infer the likely significance of results reported by others.
Second, by clearly separating prediction noise from data noise, we find that the paired prediction noise is typically greater than paired the data noise. For example, Figure 1 Right shows that the prediction noise is around two times the data noise on MATH500, which means that reducing the prediction noise can lead to a significant increase in statistical power, or a significant decrease in the minimum detectable effect size (Example 1). Even greater reductions can be expected for controlled experiments between close model pairs. When this condition holds, inferring the noise from the training curve becomes valid and useful (Example 2).
We speculate that the uniformity of these results come from the relative similarity among LLMs, since they are trained on similar data using similar methods. Like models, eval questions also have a global identity analogous to the ability to repeat experiments on the same subjects anywhere. Together, this allowed a decoupling of the noise measurements from particular experiments. Even in settings where these general patterns do not hold, our method is still applicable as long as the prediction noise is important and directly measurable.
@@
We present the main concepts informally and through examples in §2; then more precisely in §3. To ensure correctness, we validate our method by comparing with the well-established bootstrap and sign test methods (§3.5), and test our sample estimators (§3.4) against several generative models as well as on common LLM evals. §4 describes exploratory data visualization methods and the Beta model that fits the empirical distribution well. For discussions, §6.1 explains why we observe these results; §6.2 for some caveats of averaging; §6.3 on why harder questions cannot yet compensate for small sample sizes; §6.4 provide recommendations for practitioners to better deal with noise and how to use multiple evals.
The data and analysis can be found at https://all-the-noises.github.io, which shows the noise component curves, reference noise values, etc across evals, temperatures, based on millions of question-level metrics from public leaderboards and controlled experiments.
@@ 2.1 Types of noise
The prediction noise is also persistent even though we can try to reduce it by fixing the random seed, decreasing the temperature, averaging over many samples per question, or applying an aggregation method such as majority voting and verification. Fixing random seed or lowering the temperature is brittle, which only works on the exact same model. Trying to produce the best answer by aggregation is perhaps the most sound method, though this still leaves some prediction noise due to the diversity of samples. Among these noise reductions, only averaging and fixing the random seed gives the same mean while others change the mean as well. Averaging is the most general and effective, which we will focus on.
On top of sampling, additional prediction noise can come from inference-time choices such as prompt formatting, hyperparameters, and answer extraction; and training-time choices such as the random seeds, hyperparameters, and training steps. All of these can theoretically be measured by repeating the process of changing the choice, optionally training the model, and then drawing predictions on each question. Training is not typically considered noise, but it may produce more noise than signal over a small amount of steps.
@@
The eval data sampling noise comes from sampling a particular set of $N$ questions instead of another equally good $N$ questions from the unobservable population of questions. As long as the model has different expected accuracy on different questions (i.e. some questions are harder), the mean will vary when a different set of questions is sampled. Unlike the prediction noise, the data noise cannot be directly measured by evaluating on the fixed, given set of questions; and it cannot be reduced. Though the data noise can be inferred using the bootstrap or the variance estimator, it is not observed in experiments such as the training curve (Example 2) and varying the training seeds (madaan2024variance).
@@
If we only care about a specific set of questions, then we can ignore the data noise. There are real life situations when the goal is to memorize a specific list of questions such as a driving knowledge test where the set of questions is published, or to solve specific important problems (AGI/ASI, P vs. NP). On most evals, we care about the underlying ability measured by the eval rather than the specific questions of the eval, so the data noise is not to be ignored.
@@
Whereas insights, skills, and resources are required to improve LLM ability, the paired methods can increase the signal significantly. Thus they should be recommended as enormous resources are devoted to improving models. The paired method makes use of prediction differences on the same question to gain more information. To see why the paired method is better on real data, Figure 4 shows the heatmap of correctness where models at a similar overall accuracy (close in $x$-axis) also do similarly on the individual questions. Thus, the paired method has the potential to reduce the data noise significantly. Examples 2b and 1 show the potential increase in power using paired methods, especially when prediction noise > data noise, leading to the ability to detect differences several times smaller. Example 3 shows the basic idea of paired analysis.
@@
On HumanEval with $N=164$, if a model has $\bar{A}=0.5$ accuracy, then the unpaired standard error (SE) is $\sqrt{1/164\times 0.5(1-0.5)}=$ 4%. Since a factor of $\sqrt{2}$ is needed when comparing pairs of models, we need $4\%\times\sqrt{2}\times 1.96=11\%$ to get $0.05$ $p$-value. The table below shows the reductions with paired and averaging, and how the combination can significantly reduce the minimum detectable difference between typical pairs of LLMs.
@@
Unpaired with averaging 4% 8%
Paired single prediction 4% 8%
Paired with averaging 1–2% 2-4%
Table 10 of Llama 3 dubey2024llama3 reports the unpaired 95% confidence intervals of around $7\%$, thus many of the comparisons do not clear the $\sqrt{2}\times 7\%$ difference with their method.
@@
a) The training curve is invalid in theory. The training curve plots the eval result as a function of training steps and graphically shows the variance of eval results as a function of training steps (Figure 2, col 1). While separated training curves is intuitively needed for reliable results, it is not sufficient. The training curve is insufficient because it only shows the prediction noise and completely ignores the data noise as it evaluates on the same questions repeatedly. In theory, the data noise can be much higher than the prediction noise. For example, suppose the training data of models $A$ and $B$ consists of a random leak of the 100 eval questions, each included with probability 0.5. Suppose that $B$ is 5% more contaminated than $A$ just due to noise. Then $B$ has a consistently higher training curve if both $A$ and $B$ are only correct on the eval questions leaked into their training set and are wrong on the rest. Both models eventually reaching a constant peak after seeing all the leaked questions. In this example, any statistical method considering the data noise will tell us that $B$ is not better than $A$, which is true by design.
b) The training curve is valid when prediction noise is dominant. In practice, if the prediction noise dominates the data noise, then the training curve method is valid. Figure 2 shows increasingly powerful analysis on the training curve visualized by the bootstrap.
@@
When this holds, we can gain much statistical power by using the expected metric where the prediction noise is averaged away to be comparable to the data noise remaining. Figure 1 Right shows that on MATH500, the data standard error (SE) is typically less than $1/2$ of the prediction SE between different final models. Example 2b and Figure 2 show training curves evaluated on SWEBench-verified starting with the same checkpoint, the data SE is $\sim 1/6$ of the prediction SE. If we average over 36 predictions, then the total SE would be $\sqrt{2}/6=0.24$, allowing us to detect effects that are $0.24$ the size with the same $p$-value and power. Example 1 shows that the smaller reduction on HumanEval is still by a factor of 2. For controlled experiments, this allows us to detect differences several times smaller at the same significance level and power. Important caveats are discussed in §6.2.
@@
gu2024cruxeval; chiang2024chatbot tried to put error bars on a leaderboard. To do this, they use a fixed baseline for paired comparison. However, this is not generally meaningful: let $A$ be the baseline, $B$ and $B^{\prime}$ are a pair of similar predictions very different from $A$, so $\mathrm{Var}[A-B]\gg\mathrm{Var}[B-B^{\prime}]$. This example shows it’s not in general possible to have one error bar per model, or per accuracy with the paired method. However, at least for the total noise of LLM evals, this work shows that there is almost no such surprises, thus we can compute meaningful error bars.
@@ 3.2 Sampling multiple predictions on each question
We have the law of total variance satisfied by the components
@@
$\mathrm{Var}_{x,\epsilon}[A]$ is the total variance of first drawing a question $x$ from the pool of questions, and then drawing a sample answer for this question. $\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A]]$ is variance of the expected score $\mathrm{E}_{\epsilon}[A]$, which we call the data variance. $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ is the variance due to sampling from the model, which we call the prediction variance. See §A.1 for why $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ is the same as the prediction variance that can be directly measured on a fixed eval.
All noises metrics can be normalized to be their respective standard errors,
@@
3.3 Estimating from samples

### v2

@@ Abstract
Separating signal from noise is central to experiments. Applying well-established statistical methods effectively to LLM evals requires consideration of their unique noise characteristics. We clearly define and measure three types of noise: prediction noise from generating different answers on a given question, data noise from sampling questions, and their combined total noise following the law of total variance. To emphasize relative comparisons and gain statistical power, we propose the all-pairs paired method, which applies the paired analysis to all pairs of LLMs and measures all the noise components based on millions of question-level predictions across many evals and settings, revealing clear patterns. First, each eval exhibits a characteristic and highly predictable total noise level across all model pairs. Second, paired prediction noise typically exceeds paired data noise, which means reducing prediction noise by averaging can significantly increase statistical power. By measuring all the noises together, we can assess eval results in context, lowering the barrier of using the best analysis to make sound empirical decisions.
@@ 1 Introduction
Identifying signal from noise is important for interpreting experiment results. Well-established methods from statistics can increase the rigor of the conclusions drawn from data Benjamini et al. (2021). Even though these methods are in classic textbooks and clearly described for LLM evals Miller (2024), there are some choices and confusions which lead to misleading measurements and low statistical power (Example 1).
We have to consider the prediction noise of LLMs. Unlike previous probabilistic models that typically have a clear best prediction, LLMs produce persistently diverse predictions. For a model $A$ evaluated on questions $x$ with prediction seed $\epsilon$, we have
@@
The prediction variance $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ is due to the model generating different answers on a fixed question. The data variance $\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A]]$ measures how the questions vary in difficulty level and thus vary in their expected metrics. Prediction variance can be reduced but usually persists in practice. Data variance matters for reliability beyond the specific question set. The total variance is what matters in an experiment.
While prediction variance can be reduced by averaging, the data variance is irreducible and can be large when the questions vary in their difficulties. As we will see, similar models do similarly on each question, so greatly reduced data noise is possible only with the paired analysis. This combination is 3-6 times and 4 times more powerful in Example 1 and Example 2b. A difficulty is that the paired method only applies to pairs of models rather than a list of models in a table or leaderboard, thus not widely adopted. In this work, we propose the all-pairs paired method to measure all the noises, which uses the paired analysis on all pairs of results to measure the data, prediction, and total paired variances $\mathrm{Var}[A-B]$, for all model pairs $A,B$, and across many evals and settings, revealing clear patterns applicable to a table of results.
@@
For better accuracy, Figure 1 Left shows the paired total noise on SWEBench-Verified against our theoretical prediction. This allows us to have accurate-enough noise measurements without custom analysis and even to infer the likely significance of results reported by others.
Second, by clearly separating prediction noise from data noise, we find that the paired prediction noise is typically greater than the paired data noise. For example, Figure 1 Right shows that the prediction noise is around two times the data noise on MATH500, which means that reducing the prediction noise can lead to a significant increase in statistical power, or a significant decrease in the minimum detectable effect size (Example 1). Even greater reductions can be expected for controlled experiments between close model pairs. When this condition holds, inferring the noise from the training curve becomes useful (Example 2). The importance of these choices means that we must specify what noise we are measuring: data, prediction, or total noise using paired vs. unpaired method; otherwise we cannot interpret reported noise measurements Chan et al. (2025); Merrill et al. (2026).
We speculate that the uniformity of these results comes from the relative similarity among LLMs training on similar data using similar methods. Like models, eval questions also have a global identity analogous to the ability to repeat experiments on the same subjects anywhere. By summarizing the noise measurements on many pairs of results, we can put noise measurements in context and infer their likely values, lowering the barrier of the most powerful paired, averaged analysis.
@@
We present the main concepts informally and through examples in §2; then more precisely in §3. To ensure correctness, we validate our method by comparing with the well-established bootstrap and sign test methods (§E), and test our sample estimators (§A.3) against several generative models as well as on common LLM evals. §4 describes exploratory data visualization methods and the Beta model that fits the empirical distribution well. For discussions, §6.1 explains why we observe these results; §C.1 for some caveats of averaging; §6.2 on why harder questions cannot yet compensate for small sample sizes; §D provide recommendations for practitioners to better deal with noise and how to use multiple evals.
The data and analysis can be found at https://all-the-noises.github.io, which shows the results of this paper in interactive figures across evals, temperatures, based on millions of question-level metrics from public leaderboards and controlled experiments.
@@ 2.1 Types of noise
The prediction noise is also persistent even though we can try to reduce it by fixing the random seed, decreasing the temperature, averaging over many samples per question, or applying an aggregation method such as majority voting and verification. Fixing the random seed or lowering the temperature is brittle, which only works on the exact same model. Trying to produce the best answer by aggregation is perhaps the most sound method, though this still leaves some prediction noise due to the diversity of samples. Among these noise reductions, only averaging and fixing the random seed gives the same mean while others change the mean as well. Averaging is the most general and effective, which we will focus on.
Other than sampling, additional prediction noise can come from inference-time choices such as prompt formatting, hyperparameters, and answer extraction; and training-time choices such as the random seeds, hyperparameters, and training steps. All of these can in theory be measured by changing the choice then drawing predictions on each question. The prediction noise can be further decomposed to separate these effects from the model’s prediction sampling noise Romanou et al. (2026). While training is not typically considered a noise, it can result in more noise than signal over a small number of steps.
@@
The eval data sampling noise comes from sampling a particular set of $N$ questions instead of another equally good $N$ questions from the unobservable population of questions. As long as the model has different expected accuracy on different questions (i.e. some questions are harder), the mean will vary when a different set of questions is sampled. Unlike the prediction noise, the data noise cannot be directly measured by evaluating on the fixed, given set of questions; and it cannot be reduced. Though the data noise can be inferred using the bootstrap or the variance estimator, it is not observed in experiments such as the training curve (Example 2) and varying the training seeds (Madaan et al., 2024).
@@
If we only care about a specific set of questions, then we can ignore the data noise. There are real life situations when the goal is to memorize a specific list of questions such as a driving knowledge test where the set of questions is published, or to solve specific important problems (AGI/ASI, P vs. NP). On most evals, we care about the underlying ability measured by the eval rather than the specific questions of the eval, so the data noise must not be ignored.
@@
Whereas insights, skills, and resources are required to improve LLM ability, the paired methods can increase the signal significantly. Thus they should be recommended, especially given the enormous resources devoted to improving models. The paired method makes use of prediction differences on the same question to gain more information. To see why the paired method is better on real data, Figure 3 shows the heatmap of correctness where models at a similar overall accuracy (close in $x$-axis) also do similarly on the individual questions. Thus, the paired method has the potential to reduce the data noise significantly. Examples 2b and 1 show the potential increase in power using paired methods, especially when prediction noise > data noise, leading to the ability to detect differences several times smaller. Example 3 shows the basic idea of paired analysis.
@@
HumanEval Chen et al. (2021) has $N=164$ questions, if a model has $\bar{A}=0.5$ accuracy, then the unpaired standard error (SE) is $\sqrt{1/164\times 0.5(1-0.5)}=$ 4%. Since a factor of $\sqrt{2}$ is needed when comparing pairs of models, we need $4\%\times\sqrt{2}\times 1.96=11\%$ to get $0.05$ $p$-value. The table below shows the reductions with paired and averaging, and how the combination can significantly reduce the minimum detectable difference between typical pairs of LLMs.
@@
Unpaired + averaging 4% 8%
Paired 4% 8%
Paired + averaging 1–2% 2-4%
Table 10 of Llama 3 Grattafiori et al. (2024) reports the unpaired 95% confidence intervals of around $7\%$, thus many of the comparisons do not clear the $\sqrt{2}\times 7\%$ difference with their method.
@@
a) The training curve is invalid in theory. The training curve plots the eval result as a function of training steps and graphically shows the variance of eval results as a function of training steps (Figure 2, col 1). While separated training curves are intuitively needed for reliable results, it is not sufficient. The training curve is insufficient because it only shows the prediction noise and completely ignores the data noise as it evaluates on the same questions repeatedly. In theory, the data noise can be much higher than the prediction noise.
b) The training curve is valid when prediction noise is dominant. In practice, if the prediction noise dominates the data noise, then the training curve method is valid. Figure 2 shows increasingly powerful analysis on the training curve Wei et al. (2025) visualized by the bootstrap.
@@
When this holds, we can gain much statistical power by using the expected metric where the prediction noise is averaged away to be comparable to the data noise remaining. Figure 1 Right shows that on MATH500, the data standard error (SE) is typically less than $1/2$ of the prediction SE between different final models. Example 2b and Figure 2 show training curves evaluated on SWEBench-verified starting with the same checkpoint, the data SE is $\sim 1/6$ of the prediction SE. If we average over 36 predictions, then the total SE would be $\sqrt{2}/6=0.24$, allowing us to detect effects that are $0.24$ the size with the same $p$-value and power. Example 1 shows that the reduction on HumanEval is smaller but still a factor of 2. For controlled experiments, this allows us to detect differences several times smaller at the same significance level and power. Important caveats are discussed in §C.1.
@@
Gu et al. (2024); Chiang et al. (2024) tried to put error bars on a leaderboard. To do this, they use a fixed baseline for paired comparison. However, this is not generally meaningful: let $A$ be the baseline, $B$ and $B^{\prime}$ are a pair of similar predictions very different from $A$, so $\mathrm{Var}[A-B]\gg\mathrm{Var}[B-B^{\prime}]$. This example shows it’s not in general possible to have one error bar per model, or per accuracy with the paired method. However, at least for the total noise of LLM evals, this work shows that there is almost no such surprises, thus we can compute meaningful error bars.
@@ 3.2 Sampling multiple predictions on each question
While we can use $\bar{A}_{K}$ for the metric function, we decompose into components that are independent of $K$ but can be estimated using $K$ samples per question,
@@
$\mathrm{Var}_{x,\epsilon}[A]$ is the total variance of first drawing a question $x$ from the pool of questions, and then drawing a sample answer for this question. $\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A]]$ is the variance of the expected score $\mathrm{E}_{\epsilon}[A]$, which we call the data variance. $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ is the variance due to sampling from the model, which we call the prediction variance. See §A.1 for why $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A]]$ is the same as the prediction variance that can be directly measured on a fixed eval.
All the variances can be normalized to be their respective standard errors,
@@
See §A for how to estimate from samples, the importance of small-$K$ correction, and implementation details.
4 Experiments
The experimental results of this paper can be found at https://all-the-noises.github.io, where the method is tested on more evals and where the figures are interactive to support more detailed investigations and case studies. We can use this to check that the main conclusion indeed holds on all the evals. See §C.2 for a discussion of some exceptions.
4.1 Exploratory analysis and findings
4.1.1 Data heatmap
To get an overview of our data, we use a heatmap shown in Figure 3. Each row is a different question, sorted from the easiest to the hardest. Each column is a model whose $x$-coordinate is the overall accuracy $\mathrm{E}[A]$ of this model. Like models, we might consider showing questions with $y$-coordinate at their average accuracy instead of rank, but that would depend on which models are included, and less intrinsic to the eval. We immediately observe that models at a similar overall accuracy also do similarly on individual questions. Around 1000 samples are drawn for each question and model for MATH500 in Figure 3. The SWEbench figure is based on the leaderboard data, which had to commit to single answers, but the same overall pattern holds, though with more noise. On MATH500, most of the red regions of bad models are still non-zero.
4.2 All-pairs paired noises and predictable total noise
The noise components can be traded off to some extent so they can depend on the setting and not as predictable as the total. We find at more natural temperatures of $[0.7,1]$, the prediction noise tends to be higher than the data noise across evals. Figure 4 shows the noise components at two different temperatures on CRUXEval, where the breakdown differs but the total noise remains the same.
Figure 4: All noise components on CRUXEval at temperature 0.8 (left) and 0.2 (right). The prediction noise dominates at 0.8 temperature, whereas the data noise dominates at 0.2 temperature, while both still yield about the same total noise.
4.2.1 The Beta theory makes good noise predictions and fits the data
The $\operatorname{Beta}(p,1-p)$ distribution, where $p$ is the model’s mean accuracy, fits the empirical distribution of question-level expected accuracies well and predicts a paired prediction variance of p(1p)p(1-p), explaining the observed trend in Figure 1. The details are in §B.
5 Related work
Much work deals with noise in science and machine learning (Lehmann and Romano, 2005; Dror et al., 2018; Hermann et al., 2024; Card et al., 2020, among many others), we adopt the approach of Miller (2024) to estimate the variance directly. While using samples to estimate the population follows from the law of large numbers, their approach is clarifying, has unique considerations and recommendations for LLMs, and is directly generalizable to all metric functions. Their method is also consistent with the bootstrap Efron (1979) and the sign test Dixon and Mood (1946) when applicable (§E). Efron and Tibshirani (1986) clearly described estimating the standard error as the example of not needing bootstrap. The concept of total variance decomposition is present in the analysis of variance, but Section 3 of Miller (2024) applied this concept to LLMs under variance of the conditional mean and mean conditional variance, which we call the prediction noise and the data noise. They also advocated for averaging (resample), but it is not clear how much gains can be expected. In this work, we develop and test the methods of estimating noise types from samples and then empirically measure these variance components on all pairs of models.
For leaderboards, Chatbot arena Chiang et al. (2024) and CRUXEval Gu et al. (2024) computed confidence intervals using bootstrap with a fixed reference model, which allowed them to obtain per-model confidence intervals but is incorrect (Example 4). Furthermore, Chatbot arena aggregates comparisons with all other models to compute their confidence intervals, which necessarily loses any special properties of a particular pair of models.This work actually compares all pairs of models, and uses the summary statistics of all the comparisons to get meaningful noise levels.
Madaan et al. (2024) measured the noise due to random seeds, which is a form of prediction noise in our view. Any direct approach to measure the noise necessarily ignores the data noise (Example 2). For the data noise, they also proposed using unpaired Bernoulli confidence intervals. This method was adopted by Llama 3 Grattafiori et al. (2024) but is much looser than the paired methods (Example 1) and does not separate the data noise and the prediction noise. Improving model ability is hard and expensive while tightening the analysis is easy.
6 Discussions
6.1 Remarks on the main findings
Predictable total noise.
When the prediction noise is much greater than the data noise, then total noise $\approx$ prediction noise, and prediction noise is predicted by the Beta theory, so the predictable total noise is a consequence of prediction noise $\gg$ data noise. However, Figure 4 shows non-negligible data noise, yet the total noise still follows the Beta theory prediction even when the data noise exceeds the prediction noise at low temperature, except for chain-of-thought predictions. More data is needed to further test if this continues to hold, and if low temperature can reduce prediction noise as the model is training.
Prediction noise >> data noise.
Another implication is that modeling question difficulty using the Item Response Theory (IRT) or filtering questions Polo et al. (2024) should deal with the prediction noise first. Section 5 of Madaan et al. (2024) and Wang et al. (2024) find that IRT does not work well. The data filtering method $\tau$- of Wang et al. (2024) finds questions that are inversely correlated with the overall model performance. When there is high prediction noise, there is no way to get a reliable signal from a small number of questions or to reliably blame the questions for noise.
6.2 Solving hard and special problems
Leading agentic evals often have fewer than 1000 questions, with terminal bench Merrill et al. (2026) having fewer than 100 questions. In contrast, 100 multiple choice questions clearly have a lot of noise that can be realized by guessing. Perhaps the reason for believing that very small evals are okay in the agentic setting is that the setting can capture very hard problems that are not guessable. For example, if a model solved an important open problem then we would not object to the sample size of one. Here the prior probability of success is near 0, so while observing a success gives a lot of information, we are expected to fail and get no information. In other words, either we get no information or the problems are not that special. More directly, one can still use the mixed predictor (30) to generate variance from the differences in model predictions, just like guessing. Such evals still need to satisfy the regular statistical standard to be meaningful.
In human evaluations, we seem to get a lot of signal by evaluating on hard problems with a number of steps (e.g. interviews, math). The key is to reflect on the details of how the problem was solved, and if it seems to be memorized. In contrast, LLM evals and RLVR settings only give 1 bit back after doing a lot of work. If we continue in the direction of smaller and harder evals, we probably need to also generate more information per question.
6.3 Applications and Recommendations
Practical recommendations for practitioners and eval builders, including a noise-aware method for meta-analysis across multiple evals, are in §D.
6.4 Caveats of the expected metric
We recommend averaging for most controlled experiments not expected to have a large effect. However, there are caveats when targeting a particular eval metric or when the variable affects the distribution’s sharpness. See §C.1 for detailed examples.
6.5 Limitations
While the method is general, all the empirical data are from correctness evals, so whether the empirical findings generalize to other evals remains to be tested. Bowyer et al. (2025) show that CLT does not work well for less than a few hundred questions, though with 100 questions the errors are only significant in the tails. For simplicity, we did not consider the potentially impactful clustered correction of Miller (2024).
Acknowledgments.
I thank Sean O’Brien, Lovish Madaan, Dieuwke Hupkes, Weina Xie, Alex Gu, Jiawei Liu, Yuhang Lai, Linyuan Gong, and Sten Sootla for making question-level data available for analysis. I am grateful to Evan Miller, Nicolas Usunier, Zach Rait, Yuxiang Wei, Jannik Kossen, and Ari Holtzman for valuable discussions and feedback; Pedro Rodriguez, Ofir Press, Naman Jain, Baptiste Rozière, Gabriel Synnaeve, Dawn Song, and Zijian Wang for their advice and support. The all-pairs approach is inspired by Chatbot Arena and the clarity of Miller (2024) greatly helped.
Appendix A Method details
A.1 Estimating the prediction noise two ways
For a infinite set of questions, presumably $\mathrm{Var}_{\epsilon}[\mathrm{E}_{x}[A]]=0$. For $N$ specific questions, $\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A(x,\epsilon)]$ corresponds to the prediction noise that we can measure directly
$\mathrm{Var}_{\epsilon}[\mathrm{E}_{x}[A]]=\mathrm{Var}_{\epsilon}\left[\frac{1}{N}\sum_{i=1}^{N}A(x_{i},\epsilon)\right]=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}_{\epsilon}[A(x_{i},\epsilon_{i})]=\frac{1}{N}\mathrm{E}_{x}[\mathrm{Var}_{\epsilon}[A(x,\epsilon)].$ (16)
Here $\mathrm{E}_{x}[A]$ is taken over the uniform distribution of the $N$ questions. Estimating the left side directly with a finite number of random seeds $\epsilon$ has the correct expectation but has higher variance. In this case, we run the eval $K$ times and measure the variance of the $K$ final results $\mathrm{Var}[\bar{A}_{1},\ldots,\bar{A}_{K}]$ where $\bar{A}_{j}=\frac{1}{N}\sum_{i=1}^{N}A(x_{i},\epsilon_{j})$. This fails to use the fact that random samples are draw on each question independently.
If we have the ability to draw any data samples from the population, we can also measure the data noise directly.
$\mathrm{Var}_{x_{1},\ldots,x_{N}}\left[\frac{1}{N}\sum_{i=1}^{N}\mathrm{E}_{\epsilon}[A(x_{i},\epsilon)]\right]=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A(x,\epsilon)]]=\frac{1}{N}\mathrm{Var}_{x}[\mathrm{E}_{\epsilon}[A(x,\epsilon)]].$ (17)
A.2 Estimating from samples